Roberts Air Service is now in the final (terminal) year of a project.

Equipment that was purchased for the project originally cost $21 million, of which 81% has been depreciated. If Roberts can sell the used equipment for $4 million, what is the equipment's after-tax salvage value. Assume that Roberts has a corporate tax rate of 38%.

Enter your answer as a full whole dollar value (not in millions) with zero decimals. Do not enter dollar signs or commas.
Answer #1

Book value=Cost-Accumulated depreciation

=21*(1-0.81)=3.99 million

Hence gain on sale=(4-3.99)=0.01 million

Hence after-tax salvage value=Sale proceeds-(Tax rate*Gain on sale)

=4-(0.01*0.38) million

=3996200